OpenStudy (anonymous):

Simplify..... X4/5 _____ X1/3

OpenStudy (anonymous):

SHOW WORK

OpenStudy (anonymous):

\[\frac{x*\frac{4}{5}}{x*\frac{1}{3}}= \frac{\frac{4}{5}}{\frac{1}{3}} = \frac{4*3}{5*1} = \frac{12}{5}\]

OpenStudy (anonymous):

i dont understand this

OpenStudy (anonymous):

because there is an x on the top and the bottom, you can cancel it out. then you just do \[\frac{4}{5}\] divided by \[\frac{1}{3}\]

OpenStudy (anonymous):

which is 12/5

OpenStudy (anonymous):

are you positive? this is algebra 2 and we may do it a different way/method

OpenStudy (anonymous):

what is algebra 2? and did you mean, in your question \[\frac{\frac{4*x}{5}}{\frac{1*x}{3}}\] because that is what I simplified

OpenStudy (anonymous):

wht grade are you in?

OpenStudy (anonymous):

10, but I live in the UK

OpenStudy (anonymous):

algebra 2 is a high school course but im in 8th grade and im 13 and i take algebra 2

OpenStudy (turingtest):

your question is hard to decipher, is it not as 08bkrishna wrote it?